# Computing with D-Algebraic Sequences

### v1

@@ 2 Definitions
A difference-algebraic sequence over $\mathbb{K}$ is a generic zero of some non-constant difference polynomial over $\mathbb{K}$.
We view $\mathbb{K}^{\aleph_{0}}$ as the set of affine D-algebraic sequences, which we will simply refer to as D-algebraic sequences, as these are our sole focus. Consider a difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ of order $r$ and degree $d$. When $p$ is algebraically irreducible over $\mathbb{K}[x,\sigma(x),\ldots,\sigma^{r}(x)]$, its zeros are said to be defined by it. In this case, the D-algebraic sequences defined by $p$ are also said to be of order $r$ and degree $d$. Note that if $r>0$ or $d>1$, then there are infinitely many points in $\mathbb{K}^{\aleph_{0}}$ that are zeros of $p$.
@@
  • With the above definition of a difference-algebraic sequence, it is convenient to look at a difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ as the specialized (see [PSW20]) difference polynomial $p\in\mathbb{D}_{s}\coloneqq\mathbb{K}[\sigma^{j}(s(n))|j\in\mathbb{N}]$, where $s(n)$ (symbolically) represents the general term of a generic zero of $p$. This will be used to avoid lengthy notations. We will often use the automorphism $\sigma$ with the assumption that it is extended accordingly for $\mathbb{K}^{\aleph_{0}}$.
  • Let $n_{1},n_{2},n_{3}\in\mathbb{N}$, and $g$ be a bivariate rational function. To perform operations with the zeros of $p\coloneqq P(x,\sigma(x),\ldots,\sigma^{n_{1}}(x))$, $q\coloneqq Q(x,\sigma(x),\ldots,\sigma^{n_{2}}(x))\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, we write $p=P(s_{1}(n),s_{1}(n+1),\ldots,s_{1}(n+n_{1}))\in\mathbb{D}_{s_{1}}$ and $q=Q(s_{2}(n),s_{2}(n+1),\ldots,s_{2}(n+n_{2}))~{}\in~{}\mathbb{D}_{s_{2}}$, with $(s_{1}(n))_{n\in\mathbb{N}}$ and $(s_{2}(n))_{n\in\mathbb{N}}$ standing for D-algebraic sequences defined by $p$ and $q$, respectively. We aim to build $t\coloneqq T(x,\sigma(x),\ldots,\sigma^{n_{3}}(x))$ (which may be regarded as $T(s_{3}(n),s_{3}(n+1),\ldots,s_{3}(n+n_{3}))\in\mathbb{D}_{s_{3}}$) such that $g(s_{1}(n),s_{2}(n))$ is a generic zero of $t$. We refer to the difference rings $\mathbb{D}_{s_{1}},\mathbb{D}_{s_{2}},\mathbb{D}_{s_{3}}$ as the corresponding specializations of $\mathcal{D}_{\sigma}(\mathbb{K},x)$. This notation is generalized for arbitrarily many difference polynomials.
  • As the order of a difference polynomial is unchanged by application of $\sigma$, it follows that if $(s(n))_{n}$ is difference algebraic with defining difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, then any shift of $(s(n))_{n}$ is also D-algebraic and defined by $p$. The distinction between $(s(n))_{n}$ and $(s(n+i))_{n},i\in\mathbb{N}$ is apparent in their initial terms.
@@ 3 Arithmetic
We adapt the construction from [TT25, Section 2.2] to the difference case. The effectiveness of the underlying elimination theory was established in [OPS20]. The main contribution of this section is the development of systematic algorithms that automate the often tedious arithmetic of solutions to algebraic difference equations, thereby enhancing their accessibility.
@@ 3.1 Algorithm
Let $N\in\mathbb{N}\setminus\{0\}$, and consider $N$ difference-algebraic sequences as generic zeros of $N$ given difference polynomials $p_{i}\in\mathbb{D}_{s_{i}}$ of order $n_{i}\in\mathbb{N}$. Let $f\coloneqq f(s_{1},\ldots,s_{N})$, where $f\in\mathbb{K}(X_{1},\ldots,X_{N})$. We stretch that the sequence $n\mapsto f(s_{1}(n),\ldots,s_{N}(n))$ is well defined for sufficiently large integers $n$. Set $M\coloneqq\sum_{j=1}^{N}n_{i}$. We define the indeterminates
$\begin{split}&w_{j+1}(n)=s_{1}(n+j),\,\,\text{ for }\,\,0\leq j<n_{1}\\ &w_{j+1}(n)=s_{2}(n+j),\,\,\text{ for }\,\,n_{1}\leq j<n_{1}+n_{2}\\ &\vdots\\ &w_{j+1}(n)=s_{N}(n+j),\,\,\text{ for }\,\,M-n_{N}\leq j<M.\end{split}$ (4)
@@
For $i=0\ldots,N-1$, we write the difference polynomial $p_{i+1}$ as
$c_{m_{i+1}}\,{\sigma(w_{M_{i+1}})}^{m_{j}}+p_{i+1,1}\left(w_{M_{i}+1},\ldots,w_{M_{i+1}},\sigma(w_{M_{i+1}})\right),$ (6)
$c_{m_{i+1}}\in L\coloneqq\mathbb{K}[w_{M_{i}+1},\ldots,w_{M_{i+1}}]$, $p_{i+1,1}\in L[\sigma(w_{M_{i+1}})]$ such that $\deg_{\sigma(w_{M_{i+1}})}(p_{i+1,1})<m_{i+1}$. The coefficient $c_{m_{i+1}}$ is often called the initial of $p_{i+1}$. We can now view the problem as resulting from the following radical-rational dynamical system
@@
  • $E_{A_{i}}(\sigma(w_{i}))$ (with numerator $e_{a_{i}}$) is either 00 or the part of $p_{i,1}/c_{m_{i}}$ that contains $\sigma(w_{i})$ (in its numerator), $i=1,\ldots,N$;
@@
When the $N$ given difference polynomials are l.h.s., i.e., $m_{i}=1,$ $i=1,\ldots,N$, and $E_{A_{i}}=0$, $(\mathcal{M}_{f})$ is a classical rational dynamical system often considered in theoretical computer science [CDBMP23].
Let $Q$ be the least common multiple of all the denominators in the system. Without loss of generality, we assume that all $A_{i}$ and $E_{A_{i}}$ are written with $Q$ as the denominator and consider that $B=b/Q$. In the difference-algebra context, $(\mathcal{M}_{f})$ may be regarded as a system of difference polynomials on the multivariate ring of difference polynomials $\mathbb{D}_{\mathbf{w},z}\coloneqq\mathbb{D}_{w_{1},\ldots,w_{M},z}$. In this case, we consider the difference ideal
@@
where $H\coloneqq\{Q,\sigma(Q),\sigma^{2}(Q),\ldots\}$, and “$\colon H^{\infty}$” denotes the saturation with $H$.
Algorithm 1 gives the steps of our approach for the arithmetic of D-algebraic sequences.
@@
$N$ difference polynomials $p_{i}\in\mathbb{D}_{s_{i}}$ of order $n_{i}$ and a function $f\in\mathbb{K}(X_{1},\ldots,X_{n})$.
A difference polynomial of order at most $M\coloneqq n_{1}+\cdots+n_{N}$ that vanishes at $f(f_{1},\ldots,f_{N})$ for appropriate values of $n$, where each $f_{i}$ is a generic zero of $p_{i}$.
@@
Algorithm 2 Arithmetic of difference-algebraic sequences
$\mathcal{E}\subset\mathbb{K}[w_{1},\ldots,w_{M},\sigma(w_{1}),\ldots,\sigma(w_{M}),z].$
@@
    Compute the first $M-1$ shifts (application of $\sigma$) of all polynomials in $\mathcal{E}$ and add them to $\mathcal{E}$.
@@
To prove the correctness of Algorithm 1, we must show that $\langle\sigma^{\leq M}\left(I_{\mathcal{M}_{f}}\right)\rangle\cap\mathbb{K}[\sigma^{\infty}(z)]$ in step 8 is a non-trivial elimination ideal. The following theorem establishes this fact.
@@
First, let us mention that the system $(\mathcal{M}_{f})$ is consistent by Theorem 3.2 from the individual input difference polynomials. The leading monomials of
@@
Observe that $M$ is the minimal integer for which the argument of the proof of Theorem 3.2 holds.
@@ 3.2 Some closure properties and examples
Let us present some immediate consequences of the result from the previous subsection, including the ring structure of $\mathbb{K}^{\aleph_{0}}$, induced by addition and multiplication.
@@
Let $(u(n))_{n}$ and $(v(n))_{n}$ be two difference-algebraic sequences of order $r_{1}$ and $r_{2}$, respectively. Then, the following sequences are difference-algebraic of order $r$:
@@
    division: $(1/v(n))_{n\geq n_{0}}$, with $r\leq r_{2}$, provided that $(v(n))_{n}$ is non-zero and $v(n)\neq 0$ for all $n\geq n_{0}\in\mathbb{N}$;
@@
    taking radicals: ($N]{u(n)}\,)_{n},N\in\mathbb{N}$, with $r\leq r_{1}$, provided that $u(n)\geq 0$ for all $n\geq n_{0}\in\mathbb{N}$;
@@
    partial product: $(\prod_{k=k_{0}}^{n}u(k))_{n},k_{0}\in\mathbb{N}$, with $r\leq r_{1}+1$.
@@
    partial sum: $(\sum_{k=k_{0}}^{n}u(k))_{n},\,k_{0}\in\mathbb{N}$, with $r\leq r_{1}+1$;
@@
The proofs of these properties are deduced from constructions of radical-rational dynamical systems $(\mathcal{M}_{f})$ as in (7) for some rational functions, here denoted $f$.
@@
    This follows from the fact that the correctness of Algorithm 1 is unchanged with having $z$ replaced by $z^{N}$, $N\in\mathbb{N}$. Thus, $(\mathcal{M}_{f})$ is built with $f(X)=X$ such that the output equation writes $z^{N}=w_{1}$. Note, however, that this is an elementary fact. Indeed, one verifies that the difference polynomial obtained by substituting $(\sigma^{j}(x))^{i}$ by $(\sigma^{j}(x))^{i\,N}$ in $p$ has ($N]{u(n)}\,)_{n}$ as a zero.
@@
    Let us denote by $(t(n))_{n}$, the partial product of $(u(n))_{n}$. It is defined by the recursion $t(n+1)=t(n)\,u(n)$, with $t(0)=1$. First, consider the radical-rational dynamical system $(\mathcal{M}_{f})$ constructed from the difference polynomial of $(u(n))_{n}$, with $f$ unspecified. At this stage, we know that for any $f$, we have $r_{1}$ components in the system. To take $(t(n))_{n}$ into account, we add a new variable $w_{r_{1}+1}$, such that $\sigma(w_{r_{1}+1})=w_{r_{1}+1}\,w_{1}$, which represents the recursion. Then, we choose the function $b$ of the system as $b=w_{r_{1}+1}$, which tells Algorithm 1 that we want a difference polynomial for $(t(n))_{n}$. Hence, by construction, $(t(n))_{n}$ is D-algebraic of order at most $r_{1}+1$ as claimed. Some constraints of the resulting difference polynomial may define the value of $n_{0}$.
@@
    For the partial sum (also known as the partial sum), say $(s(n))_{n}$, one considers the recursion $s(n+1)=s(n)+u(n)$, with $s(0)=0$, and proceeds in the same way as with the partial product. One can verify that the difference polynomial obtained by substituting $x$ by $\sigma(x)-x$ in the defining difference polynomial of $(u(n))_{n}$ belongs to the difference ideal of the output of the algorithm. The obtained difference polynomial may need to be shifted to get its normal form.
Applications of Algorithm 1 reveal that a dynamical system in the form of $(\mathcal{M}_{f})$ from (7) serves as a certificate for computing an algebraic difference equation (or difference polynomial) associated with a given problem.
@@
    Here again, we obtain a principal ideal. The associated equation is given by {dmath} s(n) s(n+1)-s(n) s(n+2)-s(n+1)^2+s(n+1) s(n+2)+s(n+2)-s(n+3)=0.

### v2

@@ 2 Definitions
Let $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$, and $p=p_{0}p_{1}\cdots p_{\ell},$ its decomposition into irreducible components. Let $\mathbb{N}_{p_{j}}(s(n))\coloneqq\{n\in\mathbb{N},\,p_{j}(s(n))=0\}.$ A sequence $(s(n))_{n\in\mathbb{N}}$ is difference-algebraic (or D-algebraic) over $\mathbb{K}$ with defining difference polynomial $p$, if:
  1. 1.
    For all $j\in\{0,\ldots,\ell\}$, $|\mathbb{N}_{p_{j}}(s(n))|=\infty;$
  2. 2.
    For all $j\in\{0,\ldots,\ell\}$, $|\{n\in\mathbb{N}_{p_{j}}(s(n)),\,\mathcal{I}_{p_{j}}(s(n))=0\}|<\infty;$
  3. 3.
    For all $n\in\mathbb{N}$, there exists a unique $j\in\{0,\ldots,\ell\}$ such that $p_{j}(s(n))=0$.
Observe that when $\ell=0,$ Definition˜2.13 recovers Definition˜2.12: conditions $1$ and $3$ together ensure that $(s(n))_{n\in\mathbb{N}}$ is a generic zero of $p_{0}$, and condition $2$ corresponds precisely to the regularity condition. Given the “switching” behavior of a D-algebraic sequence’s terms among the irreducible factors of its defining difference polynomial (as described in Definition˜2.13), we may refer to such sequences as polymorphic D-algebraic sequences. Nevertheless, we will not explicitly distinguish between the two in this paper. From the next section onwards, our focus will primarily be on D-algebraic sequences defined by irreducible difference polynomials.
Example 2.14 (“A D-algebraic continuation of Bernoulli numbers”).
The sequence $(b(n))_{n\in\mathbb{N}}$ defined as
$\begin{split}&b\!\left(n+3\right)=-\frac{6b\!\left(n\right)b\!\left(n+1\right)+5b\!\left(n\right)b\!\left(n+2\right)-6b\!\left(n+1\right)^{2}-30b\!\left(n+1\right)b\!\left(n+2\right)}{25b\!\left(n\right)},\,n\geq 0,\\ &b\left(0\right)\coloneqq\frac{1}{2};\,b\left(1\right)\coloneqq-\frac{1}{3}\,;b\left(2\right)\coloneqq\frac{1}{6},\end{split}$ (12)
is a D-algebraic sequence whose terms are defined by the sum of two consecutive terms of a D-algebraic zero of (8). Its first three initial terms are $B_{0}+B_{1},\,B_{1}+B_{2},$ and $B_{2}+B_{3}$, obtained with the Bernoulli sequence. This shows that the sequence
$(u(n))_{n\in\mathbb{N}}=\left(1,-\frac{1}{2},\frac{1}{6},0,-\frac{1}{30},0,\frac{91}{3750},0,-\frac{423241}{11718750},0,\frac{85414689451}{915527343750},0,\ldots\right),$
defined by
$u(n)\coloneqq\begin{cases}b(n)\quad\text{if}\quad n=2k,\,k\geq 1,\\ B_{n}\quad\text{ otherwise, }\end{cases}$ (13)
is a generic zero of (8), which, as the Bernoulli sequence, cannot be defined by it. By adjusting (12), we deduce the following difference polynomial to define $(u(n))_{n\in\mathbb{N}}$ as a D-algebraic sequence.
$\sigma^{2}(x)\,\left(25\,x\sigma^{4}(x)+11\,x\sigma^{2}(x)-36\,(\sigma^{2}(x))^{2}+\sigma^{3}(x)\right).$ (14)
For $n\geq 0$, the first factor from the left helps in computing zero terms at odd indices, and the second factor helps to compute the nonzero terms. $\blacksquare$
Example 2.15.
Consider $p\coloneqq\left(2x\sigma(x)+x+\sigma(x)\right)\left(2\sigma(x)x-x-\sigma(x)\right)$. Using $p$ and the initial term $s(0)\coloneqq 1$, we define the recursion:
$s(n+1)\coloneqq\begin{cases}\frac{s\left(n\right)}{2\cdot s\left(n\right)-1},\,\text{if}\,\,\,n\,\,\,\text{is odd,}\\ -\frac{s\left(n\right)}{2\cdot s\left(n\right)+1},\,\text{if}\,\,\,n\,\,\,\text{is even.}\end{cases}$ (15)
The sequence $(s(n))_{n\in\mathbb{N}}$ is D-algebraic and we have $s(n)=\frac{(-1)^{n}}{2n+1},$ for all $n\in\mathbb{N}$. This can be shown by checking the initial term and substituting $\frac{(-1)^{n}}{2n+1}$ for $n=2k$ and $n=2k+1$ into (15) to verify that $s(n+1)=\frac{(-1)^{n+1}}{2(n+1)+1}.$ Using the algorithm from [TTW24], one shows that $(s(n))_{n\in\mathbb{N}}$ can also be defined with the difference polynomial $q\coloneqq(2\,x+\sigma(x))\,\sigma^{2}(x)+x\sigma(x).$ A question of interest for future studies arose in the understanding of the relation between $p$ and $q$ for $(s(n))_{n\in\mathbb{N}}$. $\blacksquare$
We view $\mathbb{K}^{\aleph_{0}}$ as the set of affine D-algebraic sequences, which we will simply refer to as D-algebraic sequences, as these are our sole focus. Although arithmetic with coordinates does not have a particular interest in algebraic geometry, we anticipate a comparative study relating projective D-algebraic sequences, which we may write $\mathbb{P}^{\aleph_{0}}(\mathbb{K})$, and $\mathbb{K}^{\aleph_{0}}$, especially on their defining difference polynomials.
@@
With the above definition of a difference-algebraic sequence, it is convenient to look at an irreducible difference polynomial $p\in\mathcal{D}_{\sigma}(\mathbb{K},x)$ as the specialized (see [PSW20]) difference polynomial $p\in\mathbb{D}_{s}\coloneqq\mathbb{K}[\sigma^{j}(s(n))|j\in\mathbb{N}]$, where $s(n)$ (symbolically) represents the general term of a D-algebraic zero of $p$. This will be used to avoid lengthy notations. We use the automorphism $\sigma$ with the assumption that it is extended accordingly for $\mathbb{K}^{\aleph_{0}}$.
@@ 3 Arithmetic
We adapt the construction from [TT25, Section 2.2] to the difference case. We focus on irreducible difference polynomials.
Let $N\in\mathbb{N}\setminus\{0\}$. We consider $N$ D-algebraic sequences $(s_{i}(n))_{n\in\mathbb{N}},$ each defined by a given irreducible difference polynomial $p_{i}\in\mathcal{D}_{\sigma}(\mathbb{K},x),$ of order $n_{i}\in\mathbb{N}.$ Let $g,h\in\mathbb{K}[X_{1},\ldots,X_{N}]$ be two coprime polynomials, and define $f\coloneqq\frac{g}{h}$. We aim to construct a difference polynomial $q$ such that the sequence defined by $t(n)=f(s_{1}(n),\ldots,s_{N}(n))$ is a D-algebraic zero of $q.$
For $t(n)$ to be well-defined for sufficiently large $n$, the denominator $h(s_{1}(n),\ldots,s_{N}(n))\neq 0$ must be nonzero. Combining this requirement with the regularity condition for each individual sequence $(s_{i}(n))_{n\in\mathbb{N}}$, we deduce the following critical prerequisite for the construction:
$Q(s_{1}(n),\ldots,s_{N}(n))\coloneqq h(s_{1}(n),\ldots,s_{N}(n))\,\prod_{i=1}^{N}\mathcal{I}_{p_{i}}(s_{i}(n))\,\neq\,0,\,\,\text{for sufficiently large}\,\,n.$ (16)
For our algorithmic approach, we utilize the symbolic representation $p_{i}(s_{i}(n),\ldots,s_{i}(n+n_{i}))\in\mathbb{D}_{s_{i}}.$ This signifies that we are treating the $s_{i}(n)$’s as symbolic variables, rather than focusing on the specific numerical values dependent on $n$. We use these $p_{i}$’s and $f$, combined with condition (16), to construct a difference polynomial $q(t(n),\ldots,t(n+n_{q}))\in\mathbb{D}_{t},n_{q}\in\mathbb{N}.$ This $q$ is designed to vanish when evaluated at $t(n)=f(s_{1}(n),\ldots,s_{N}(n))$, provided that relevant shifts of $Q(s_{1}(n),\ldots,s_{N}(n))$ are nonzero. In practice, due to the D-algebraic nature of the sequences $s_{i}$’s, this condition simplifies to assuming that corresponding shifts of $h(s_{1}(n),\ldots,s_{N}(n))$ are nonzero.
While precisely determining the indices $n$ for which $Q(s_{1}(n),\ldots,s_{N}(n))$ might vanish is generally challenging, this underscores the importance and utility of an algorithmic method that can derive $q$ under the assumption of condition (16)’s general validity. We also mention that the initial terms of the sequences play a crucial role in the practical application of our method.
@@ 3.1 Algorithm
We use the same notations from the introductory paragraph of this section. We define the indeterminates
[…]
@@
For $i=0\ldots,N-1$, we write $c_{m_{i}+1}=\mathcal{I}_{p_{i+1}}$, and $p_{i+1}$ as (see (10))
$c_{m_{i+1}}\,{\sigma(w_{M_{i+1}})}^{m_{j}}+p_{i+1,1}\left(w_{M_{i}+1},\ldots,w_{M_{i+1}},\sigma(w_{M_{i+1}})\right).$ (19)
We can now view the problem as resulting from the following radical-rational dynamical system
@@
  • $E_{A_{i}}(\sigma(w_{i}))$ (with numerator $e_{a_{i}}$) is either 0 or the part of $p_{i,1}/c_{m_{i}}$ that contains $\sigma(w_{i})$ (in its numerator), $i=1,\ldots,N$;
@@
When the $N$ given difference polynomials are rationalizing, i.e., $m_{i}=1,$ $i=1,\ldots,N$, and $E_{A_{i}}=0$, $(\mathcal{M}_{f})$ is a classical rational dynamical system often considered in theoretical computer science [CDBMP23].
Let $Q$ be the common denominator of all denominators in the system. Without loss of generality, we assume that all $A_{i}$ and $E_{A_{i}}$ are written with $Q$ as the denominator and consider that $B=b/Q$. This $Q$ carries the condition from (16). We may regard $(\mathcal{M}_{f})$ as a system of difference polynomials on the multivariate ring of difference polynomials $\mathbb{D}_{\mathbf{w},z}\coloneqq\mathbb{D}_{w_{1},\ldots,w_{M},z}$. The corresponding difference ideal is
@@
where $H\coloneqq\{Q,\sigma(Q),\sigma^{2}(Q),\ldots\}$, and $\colon H^{\infty}$” denotes the saturation with $H$.
Algorithm˜1 gives the steps of our approach for the arithmetic of D-algebraic sequences.
@@
NN difference polynomials $p_{i}\in\mathbb{D}_{s_{i}}$ of order $n_{i}$ and a function $f\in\mathbb{K}(X_{1},\ldots,X_{n})$.
A difference polynomial of order at most $M\coloneqq n_{1}+\cdots+n_{N}$ that vanishes at $f(s_{1},\ldots,s_{N})$ for appropriate values of $n$, where each $s_{i}$ is a D-algebraic zero of $p_{i}$.
@@
    Compute the first M1M-1 shifts (application of $\sigma$) of all polynomials in $\mathcal{E}$ and add them to $\mathcal{E}$.
@@
To prove the correctness of Algorithm˜1, we must show that $\langle\sigma^{\leq M}\left(I_{\mathcal{M}_{f}}\right)\rangle\cap\mathbb{K}[\sigma^{\infty}(z)]$ in step 8 is a non-trivial elimination ideal. The following theorem establishes this fact.
@@
First, let us mention that the system $(\mathcal{M}_{f})$ is consistent by Theorem˜3.2 from the individual input difference polynomials. The leading monomials of
@@
Observe that $M$ is the minimal integer for which the argument of the proof of Theorem˜3.2 holds.
@@ 3.2 Some closure properties and examples
We present some immediate consequences of the result from the previous subsection, including the ring structure of $\mathbb{K}^{\aleph_{0}}$, induced by addition and multiplication. We show how to use Algorithm˜1 beyond arithmetic properties. The central fact resulting from Theorem˜3.2 is that building a dynamical system in the form of $\mathcal{M}_{f}$ in (20) is enough to show the existence of a difference polynomial vanishing at sequences’ transformations encoded by some $f$. We exploit the fact that the form of the system is unchanged by adding new variables and raising them to some powers on its left-hand side.
We say that a sequence $(s(n))_{n\in\mathbb{N}}$ (or simply $(s(n))_{n}$ if there is no ambiguity) is D-algebraic of order $r$ if it is defined by an irreducible difference polynomial of order $r$.
@@
Let $(u(n))_{n}$ and $(v(n))_{n}$ be two difference-algebraic sequences of order $r_{1}$ and $r_{2}$, respectively, and $i,j\in\mathbb{N},i<r_{1},\,j<r_{2}$. Then, the following sequences are difference-algebraic of order $r$:
@@
    division: $(1/v(n+j))_{n\geq n_{0}}$, with $r\leq r_{2}$, $v(n)\neq 0$ for all $n\geq n_{0}\in\mathbb{N}$;
@@
    taking radicals: $(\sqrt[N]{u(n+i)}\,)_{n},N\in\mathbb{N}$, with $r\leq r_{1}$;
@@
    partial product: $(\prod_{k=k_{0}}^{n}u(k+i))_{n},k_{0}\in\mathbb{N}$, with $r\leq r_{1}+1;$
@@
    partial sum: $(\sum_{k=k_{0}}^{n}u(k+i))_{n},\,k_{0}\in\mathbb{N}$, with $r\leq r_{1}+1$;
@@
The proofs of these properties are deduced from constructions of radical-rational dynamical systems $(\mathcal{M}_{f})$ as in (20) for some rational functions, here denoted $f$. We provide details for $i=j=0$ as the case $i,j\neq 0$ only implies a different encoding of $f$ with the variables of the dynamical system.
@@
    This follows from the fact that the correctness of Algorithm˜1 is unchanged with having $z$ replaced by $z^{N}$, $N\in\mathbb{N}$. Thus, $(\mathcal{M}_{f})$ is built with $f(X)=X$ such that the output equation writes $z^{N}=w_{1}$. Note, however, that this is an elementary fact. Indeed, one verifies that the difference polynomial obtained by substituting $(\sigma^{j}(x))^{i}$ by $(\sigma^{j}(x))^{i\,N}$ in $p$ has $(\sqrt[N]{u(n)}\,)_{n}$ as a zero.
@@
    Let us denote by $(t(n))_{n}$, the partial product of $(u(n))_{n}$. It is defined by the recursion t(n+1)=t(n)u(n)t(n+1)=t(n)\,u(n), with $t(0)=1$. First, consider the radical-rational dynamical system $(\mathcal{M}_{f})$ constructed from the difference polynomial of $(u(n))_{n}$, with $f$ unspecified. At this stage, we know that for any $f$, we have $r_{1}$ components in the system. To take $(t(n))_{n}$ into account, we add a new variable $w_{r_{1}+1}$, such that $\sigma(w_{r_{1}+1})=w_{r_{1}+1}\,w_{1}$, which represents the recursion. Then, we choose the function $b$ of the system as $b=w_{r_{1}+1}$, which tells Algorithm˜1 that we want a difference polynomial for $(t(n))_{n}$. Hence, by construction, $(t(n))_{n}$ is D-algebraic of order at most $r_{1}+1$ as claimed. Some constraints of the resulting difference polynomial may define the value of $k_{0}$.
@@
    For the partial sum, say $(s(n))_{n}$, one considers the recursion $s(n+1)=s(n)+u(n)$, with $s(0)=0$, and proceeds in the same way as with the partial product. One can verify that the difference polynomial obtained by substituting $x$ by $\sigma(x)-x$ in the defining difference polynomial of $(u(n))_{n}$ belongs to the difference ideal of the output of the algorithm.
Applications of Algorithm˜1 reveal that a dynamical system in the form of $(\mathcal{M}_{f})$ from (20) serves as a certificate for computing an algebraic difference equation (or difference polynomial) associated with a given problem.
In the examples below, the assumption related to (16) is easily verified. Therefore, we will not pay particular attention to that required condition.
@@
    Here again, we obtain a principal ideal. The associated equation is given by
    $s(n)s(n+1)-s(n)s(n+2)-s(n+1)^{2}+s(n+1)s(n+2)+s(n+2)-s(n+3)=0.$ (28)